OpenStudy (anonymous):

What is the solution of log3x - 12 729 = 6? the 3x-12 is like lower than log

OpenStudy (jdoe0001):

\(\huge log_{3x - 12}(729) = 6\quad ?\)

OpenStudy (anonymous):

yes like that

OpenStudy (ranga):

\[\Large \log_{3x - 12} 729 = 6\]\[\Large \log_{B} A = C\]means\[\Large A = B^{C}\]

OpenStudy (jdoe0001):

hmm you know the change of base rule, right?

OpenStudy (ranga):

\[\Large \log_{3x - 12}(729) = 6\quad means \quad (3x - 12)^{6} = 729\]

OpenStudy (ranga):

729 = 3^6 Solve for x.

OpenStudy (anonymous):

uhm so like distribute th e6 to the 3x and the 12?

OpenStudy (ranga):

\[\Large (3x - 12)^{6} = 729 = 3^{6}\]. Both are raised to 6. Therefore, 3x - 12 = 3. Solve for x.

OpenStudy (anonymous):

so it would be 5?

OpenStudy (ranga):

Yes, x = 5. Another way to look at it is: (3x - 12)^6 = 729 Find the sixth root on both sides (that is, raise both sides to the power 1/6): 3x - 12 = 729^(1/6) = 3